# Homological stratification and descent

### v1

@@ Abstract.
We introduce a notion of stratification for big tensor-triangulated categories relative to the homological spectrum and develop the fundamental features of this theory. In particular, we demonstrate that it exhibits excellent descent properties. In conjunction with Balmer’s Nerves of Steel conjecture, we conclude that stratification admits a general form of descent. This gives a uniform treatment of several recent stratification results and provides a complete answer to the question: When does stratification descend? As a new application, we extend earlier work on the tensor triangular geometry of equivariant module spectra from finite groups to compact Lie groups.
@@ Introduction
In practice, $R$ is often taken to be the (graded) endomorphism ring $[\mathbb{1},\mathbb{1}]_{*}$ of the unit in $\mathscr{T}$ [^2] More generally, cohomological stratification operates in the setting of arbitrary compactly generated triangulated categories, i.e., it does not require a tensor structure on $\mathscr{T}$. A prominent application of this theory is the stratification of the stable module category of a finite group in modular characteristic due to Benson, Iyengar and Krause [BIK11a]. Key to their proof is a descent result for cohomological stratification that provides a reduction from finite groups to elementary abelian groups, which can then be tackled via Neeman’s theorem.
A significant drawback of cohomological stratification is that it can only apply in situations where the parametrizing object is affine and noetherian. Addressing both issues simultaneously, in [BHS23b] the three first-named authors introduced a notion of tensor-triangular stratification which is instead based on the Balmer–Favi notion of support [BF11]. The latter takes values in $\operatorname{Spc}(\mathscr{T}^{c})$ and tensor-triangular stratification asserts that it provides a bijection
[…]
@@
$\rho\colon\operatorname{Spc}(\mathscr{T}^{c})\to\operatorname{Spec}([\mathbb{1},\mathbb{1}]_{*})$
and applies to numerous new examples. However, two serious issues remain: Firstly, it still requires a finiteness assumption on the spectrum through the topological hypothesis of being “weakly noetherian”. Secondly, while the permanence of stratification has been established via many instances of descent (e.g., Zariski descent, finite étale descent, nil descent, etc; see [BHS23b, BCHS23, BCH+24b]), a general descent statement remains elusive.
@@
  • (Generality) Homological stratification works without any point-set topological restrictions on the spectrum, thus avoiding the finiteness assumptions required for tensor-triangular and cohomological stratification.
@@
homological stratification tensor-triangular stratification cohomological stratification ${\operatorname{Spc}^{h}(\mathscr{T}^{c})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${\operatorname{Spc}(\mathscr{T}^{c})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho}$${\operatorname{Spec}([\mathbb{1},\mathbb{1}]_{*})}$
@@
Here $E_{\mathscr{B}}$ is a pure-injective object in $\mathscr{T}$ associated to each $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$; see 1.1. The equivalence of the first two conditions is completely analogous to the characterization of cohomological stratification and tt-stratification established in [BIK11b, Theorem 4.2] and [BHS23b, Theorem 4.1], respectively. The third characterization is new and involves a notion of homological cosupport $\operatorname{Cosupp}^{h}$ which we introduce and study in this paper, motivated by its tt-theoretic counterpart studied in [BCHS23].
@@
This result generalizes the étale descent theorem from [BHS23b, Theorem 6.4] and [Bar22, Section 2.2.2] by removing the separability condition present there. As a special case, we obtain the following result, which generalizes the descent theorem of [BCH+24b, Example 12.18]:
@@
We also generalize nil-descent and quasi-finite descent established in [BCHS23] to families of geometric functors:
Theorem H (10.1$(a)(b)$).
@@
This extends [BHS23b, Theorem 15.1] from finite groups to compact Lie groups. Indeed, there we proved the finite analog of I under the assumption that $\operatorname{Mod}(R)$ was tt-stratified, which implies both the Nerves of Steel conjecture and h-stratification by E.
@@
We compare the homological support with the Balmer–Favi support in Section 5 and establish several base change results for the homological support in Section 6. Of particular note is 6.4, which establishes that the homological support satisfies the Avrunin–Scott identity whenever the h-detection property holds. In Section 7, we introduce the notion of a weakly descendable family of functors and argue that this is the appropriate context in which to study descent. This culminates in 7.21. We then turn to the comparison between h-stratification and tt-stratification in Section 8 yielding 8.6 and its corollaries. Then in Section 9 we turn to the task of descending the Nerves of Steel conjecture, yielding 9.2 and 9.4. We also show that the Nerves of Steel conjecture is equivalent to a purely support-theoretic statement (9.6). Our descent results are then applied in Section 10 to obtain several specific descent results for tt-stratification, including 10.3 among others. Applications to equivariant spectra are then given in Section 11, including 11.9. Finally, we conclude the paper in Section 12 with a list of open problems.
@@
  • Although not necessary to understand the logic of the results, some familiarity with the tt-theory of stratification and costratification developed in [BHS23b] and [BCHS23] and the descent results of [BCH+24b] will be helpful to contextualize the results of this paper.
@@ Acknowledgments
We thank Scott Balchin and Paul Balmer for helpful discussions. TB is supported by the European Research Council (ERC) under Horizon Europe (grant No. 101042990) and would like to thank the Max Planck Institute for its hospitality. DH is supported by grant number TMS2020TMT02 from the Trond Mohn Foundation.
@@ 1. Homological support and cosupport
The homological cosupport does not appear to have been previously studied in this level of generality, but it will play its own role in the story, as in [BCHS23].
@@
The ()($\Rightarrow$)( ⇒ ) direction holds unconditionally by 1.4, so it suffices to prove the ()($\Leftarrow$)( ⇐ ) implication. Let $t\in\mathscr{T}$ with $t\otimes E_{\mathscr{B}}\neq 0$. By our minimality assumption, this implies that $E_{\mathscr{B}}\in\operatorname{Locid}\langle t\otimes E_{\mathscr{B}}\rangle$. If we had $\mathsf{hom}(t\otimes E_{\mathscr{B}},E_{\mathscr{B}})=0$, then $\mathsf{hom}(E_{\mathscr{B}},E_{\mathscr{B}})=0$ which contradicts $E_{\mathscr{B}}\neq 0$. Therefore, we see
@@ 2. The homological local-to-global principle
The equivalence of $(a)$ and $(d)$ follows verbatim from the argument in [BCHS23, Theorem 6.4] which establishes the equivalence of the classical LGP property and the classical codetection property. On the other hand, part $(a)$ is equivalent to part $(b)$ by [BCHS23, (2.6)]. Finally, the equivalence of $(b)$ and $(c)$ is immediate from the definition of the naive h-support. ∎
@@
The h-LGP property of 2.1 implies the equivalent conditions of 2.2. However, a priori the latter conditions are weaker since we do not know whether $\operatorname{Supp}^{h}(t)=\operatorname{Supp}^{n}(t)$ in general. In particular, it is not immediate from the above whether the h-codetection property implies the h-detection property. This is in contrast to the situation in [BCHS23]. However, we have the following conditional result:
@@
Suppose that $(a)$ holds, so that in particular h-detection holds. It follows from 1.10 that $(b)$ holds. Conversely, the equality in $(b)$ implies that
@@
so the h-LGP follows from h-codetection by 2.2. ∎
@@
Note that the homological local-to-global principle implies the homological detection property (1.9). The following example shows that the converse does not hold.
@@ 3. Homological stratification
We now introduce homological stratification and establish its basic properties. We begin with the following homological variant of [BHS23b, Theorem 4.1].
@@
by [BCHS23, (13.4)]. This proves that $\operatorname{Locid}\langle f_{*}(\mathbb{1})\rangle$ is minimal. ∎
@@
Let $\mathscr{T}=\operatorname{SH}$ be the stable homotopy category. By [Bal20b, Corollary 5.10], which relies on the nilpotence theorem of Devinatz–Hopkins–Smith [DHS88], the canonical map $\operatorname{Spc}^{h}(\operatorname{SH}^{c})\to\operatorname{Spc}(\operatorname{SH}^{c})$ is a bijection; cf. 8.1 below. In particular, for a given prime number $p$, there is a homological prime $\mathscr{B}_{p,\infty}$ corresponding to the tt-prime of all finite $p$-torsion spectra. The associated weak ring identifies with the mod $p$ Eilenberg–Mac Lane spectrum $H{\mathbb{F}_{p}}$, i.e., $E_{\mathscr{B}_{p,\infty}}=H{\mathbb{F}_{p}}$; see [BC21, Corollary 3.6]. However, it follows from Ravenel’s work [Rav84] that there are strict inclusions
$0\subsetneq\operatorname{Locid}\langle I\rangle\subsetneq\operatorname{Locid}\langle H{\mathbb{F}_{p}}\rangle$
where $I$ denotes the $p$-local Brown–Comenetz dual of the sphere spectrum. Indeed, on the one hand, we have $I\in\operatorname{Locid}\langle H{\mathbb{F}_{p}}\rangle$ because the homotopy groups of $I$ are $p$-torsion and concentrated in non-negative degrees. On the other hand, we have $H{\mathbb{F}_{p}}\not\in\operatorname{Locid}\langle I\rangle$ because $H{\mathbb{F}_{p}}\otimes H{\mathbb{F}_{p}}\neq 0$ while $H{\mathbb{F}_{p}}\otimes I=0$.
@@
The spectrum in 3.14 is a single point, which shows that there is no obvious condition on the space $\operatorname{Spc}^{h}(\mathscr{T}^{c})$ that ensures that the homological local-to-global principle holds. This should be contrasted with the classical local-to-global principle, which holds whenever $\operatorname{Spc}(\mathscr{T}^{c}$) is noetherian [BHS23b, Theorem 3.22].
@@
Recall that the Bousfield class $A(t)$ of an object $t\in\mathscr{T}$ is the localizing ideal $A(t)\coloneqq\ker(-\otimes t)$ and that the Bousfield lattice of $\mathscr{T}$ is the lattice of Bousfield classes ordered by reverse inclusion; see [BHS23b, Section 8]. Surprisingly, the following proposition shows that the classification of Bousfield classes is equivalent to the h-detection property whenever the homological support agrees with the naive homological support. Indeed, it is immediate from the definition of naive homological support that $\operatorname{Supp}^{n}(t)\subseteq\operatorname{Supp}^{n}(s)$ whenever $A(t)\leq A(s)$. Therefore, there is a well-defined order-preserving map
[…]
@@
For any $\mathscr{T}$ satisfying the h-detection property, we obtain a description of the meet operation on its Bousfield lattice: $A(s)\wedge A(t)=A(s\otimes t)$; cf. [BHS23b, Theorem 8.8].
@@
If $\mathscr{T}$ is h-stratified then every localizing ideal is a Bousfield class and the lattice of localizing ideals of $\mathscr{T}$ is isomorphic to the Bousfield lattice of $\mathscr{T}$; cf. [BHS23b, Theorem 8.8].
@@ 4. Homological stratification via cosupport
We now establish characterizations of homological stratification in terms of the behaviour of homological cosupport. The proof of the next proposition closely follows its tensor-triangular counterpart ([BCHS23, Theorem 7.15]):
@@
so $\mathsf{hom}(t_{1},t_{2})\neq 0$ by $(c)$. The minimality criterion [BCHS23, Lemma 7.12] then shows that $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle$ is minimal. ∎
@@
4.2 is a homological analogue of the characterization of stratification in terms of cosupport [BCHS23, Theorem 7.15]. The goal of the rest of this section is to establish a variant of this result formulated purely in terms of support and cosupport.
@@
The proof of the next theorem makes use of Brown–Comenetz duality. Recall from [BCHS23, Definition 9.4] that the Brown–Comenetz dual $I_{c}$ of a compact object $c\in\mathscr{T}$ is characterized by the formula
@@
for all $t\in\mathscr{T}$. Following the proof of [BCHS23, Proposition 12.9], we compute

### v2

@@ Abstract.
We introduce a notion of stratification for rigidly-compactly generated tensor-triangulated categories relative to the homological spectrum and develop the fundamental features of this theory. In particular, we demonstrate that it exhibits excellent descent properties. In conjunction with Balmer’s Nerves of Steel conjecture, we conclude that classical stratification also admits a general form of descent. This gives a uniform treatment of several recent stratification results and provides a complete answer to the question: When does stratification descend? As a new application, we extend earlier work on the tensor triangular geometry of equivariant module spectra from finite groups to compact Lie groups.
@@ Introduction
In practice, $R$ is often taken to be the (graded) endomorphism ring $\operatorname{End}_{\mathscr{T}}^{\bullet}(\mathbb{1})$ of the unit in $\mathscr{T}$. [^2] More generally, cohomological stratification operates in the setting of arbitrary compactly generated triangulated categories, i.e., it does not require a tensor structure on $\mathscr{T}$. A prominent application of this theory is the stratification of the stable module category of a finite group in modular characteristic due to Benson, Iyengar and Krause [BIK11a]. Key to their proof is a descent result for cohomological stratification that provides a reduction from finite groups to elementary abelian groups, which can then be tackled via Neeman’s theorem.
A significant drawback of cohomological stratification is that it can only apply in situations where the parametrizing object is affine and noetherian. Addressing both issues simultaneously, in [BHS23d] the three first-named authors introduced a notion of tensor-triangular stratification which is instead based on the Balmer–Favi notion of support [BF11]. The latter takes values in $\operatorname{Spc}(\mathscr{T}^{c})$ and tensor-triangular stratification asserts that it provides a bijection
$\operatorname{Supp}\colon\big\{\text{localizing ideals of $\mathscr{T}$}\big\}\xrightarrow{\cong}\big\{\text{subsets of $\operatorname{Spc}(\mathscr{T}^{c})$}\big\}.$
@@
$\rho\colon\operatorname{Spc}(\mathscr{T}^{c})\to\operatorname{Spec}(\operatorname{End}_{\mathscr{T}}^{\bullet}(\mathbb{1}))$
and applies to numerous new examples. However, two serious issues remain: Firstly, it still requires a topological assumption on the spectrum, namely the hypothesis of being “weakly noetherian”. Secondly, while the permanence of stratification has been established via many instances of descent (e.g., Zariski descent, finite étale descent, nil descent, etc; see [BHS23d, BCH+23b, BCH+24a]), a general descent statement remains elusive.
@@
  • (Generality) Homological stratification works without any point-set topological restrictions on the spectrum, thus avoiding the (weakly) noetherian assumptions required for tensor-triangular and cohomological stratification.
@@
homological stratification tensor-triangular stratification cohomological stratification ${\operatorname{Spc}^{h}(\mathscr{T}^{c})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${\operatorname{Spc}(\mathscr{T}^{c})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho}$${\operatorname{Spec}(\operatorname{End}_{\mathscr{T}}^{\bullet}(\mathbb{1}))}$
@@
Here $E_{\mathscr{B}}$ is a pure-injective object in $\mathscr{T}$ associated to each $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$; see 1.1. The equivalence of the first two conditions is completely analogous to the characterization of cohomological stratification and tt-stratification established in [BIK11b, Theorem 4.2] and [BHS23d, Theorem 4.1], respectively. The third characterization is new and involves a notion of homological cosupport $\operatorname{Cosupp}^{h}$ which we introduce and study in this paper, motivated by its tt-theoretic counterpart studied in [BCH+23b].
@@
This result generalizes the étale descent theorem from [BHS23d, Theorem 6.4] and [BAR22, Section 2.2.2] by removing the separability condition present there. As a special case, we obtain the following result, which generalizes the descent theorem of [BCH+24a, Example 12.18]:
@@
We also generalize nil-descent and quasi-finite descent established in [BCH+23b] to families of geometric functors:
Theorem H (10.1(a)(b)(a)(b)).
@@
This extends [BHS23d, Theorem 15.1] from finite groups to compact Lie groups. Indeed, there we proved the finite analog of I under the assumption that $\operatorname{Mod}(R)$ was $t$-stratified, which implies both the Nerves of Steel conjecture and h-stratification by E.
@@
We compare the homological support with the Balmer–Favi support in Section 5 and establish several base change results for the homological support in Section 6. Of particular note is 6.4, which establishes that the homological support satisfies the Avrunin–Scott identity whenever the h-detection property holds. In Section 7, we introduce the notion of a weakly descendable family of functors and argue that this is the appropriate context in which to study descent. This culminates in 7.22. We then turn to the comparison between h-stratification and tt-stratification in Section 8 yielding 8.6 and its corollaries. Then in Section 9 we turn to the task of descending the Nerves of Steel conjecture, yielding 9.2 and 9.4. We also show that the Nerves of Steel conjecture is equivalent to a purely support-theoretic statement (9.6). Our descent results are then applied in Section 10 to obtain several specific descent results for tt-stratification, including 10.3 among others. Applications to equivariant spectra are then given in Section 11, including 11.9. Finally, we conclude the paper in Section 12 with a list of open problems.
@@
  • Although not necessary to understand the logic of the results, some familiarity with the $t$-theory of stratification and costratification developed in [BHS23d] and [BCH+23b] and the descent results of [BCH+24a] will be helpful to contextualize the results of this paper.
@@ Acknowledgments
We thank Scott Balchin and Paul Balmer for helpful discussions, as well as Jordan Williamson and the anonymous referee for useful comments on an earlier version. TB is supported by the European Research Council (ERC) under Horizon Europe (grant No. 101042990) and would like to thank the Max Planck Institute for its hospitality. DH is supported by grant number TMS2020TMT02 from the Trond Mohn Foundation.
@@ 1. Homological support and cosupport
The homological cosupport does not appear to have been previously studied in this level of generality, but it will play its own role in the story, as in [BCH+23b].
@@
The $(\Rightarrow)$ direction holds unconditionally by 1.4, so it suffices to prove the $(\Leftarrow)$ implication. Let $t\in\mathscr{T}$ with $t\otimes E_{\mathscr{B}}\neq 0$. By our minimality assumption, this implies that $E_{\mathscr{B}}\in\operatorname{Locid}\langle t\otimes E_{\mathscr{B}}\rangle$. If we had $\mathsf{hom}(t\otimes E_{\mathscr{B}},E_{\mathscr{B}})=0$, then $\mathsf{hom}(E_{\mathscr{B}},E_{\mathscr{B}})=0$ which contradicts $E_{\mathscr{B}}\neq 0$. Therefore, we see
@@ 2. The homological local-to-global principle
The equivalence of $(a)$ and $(d)$ follows verbatim from the argument in [BCH+23b, Theorem 6.4] which establishes the equivalence of the classical LGP property and the classical codetection property. On the other hand, part $(a)$ is equivalent to part $(b)$ by [BCH+23b, (2.6)]. Finally, the equivalence of $(b)$ and $(c)$ is immediate from the definition of the naive h-support. ∎
@@
The h-LGP property of 2.1 implies the equivalent conditions of 2.2. However, a priori the latter conditions are weaker since we do not know whether $\operatorname{Supp}^{h}(t)=\operatorname{Supp}^{n}(t)$ in general. In particular, it is not immediate from the above whether the h-codetection property implies the h-detection property. This is in contrast to the situation in [BCH+23b]. However, we have the following conditional result:
@@
Suppose that $(a)$ holds, so that in particular h-detection holds. Indeed, $\operatorname{Supp}^{h}(t)=\varnothing$ implies $\operatorname{Locid}\langle t\rangle=0$ by the h-LGP, hence $t=0$. It then follows from 1.10 that $(b)$ holds. Conversely, the equality in $(b)$ implies that
@@
so the h-LGP follows from h-codetection by $2.$ ∎
@@
We remind the reader that the homological local-to-global principle implies the homological detection property (1.9). The following example shows that the converse does not hold.
@@ 3. Homological stratification
We now introduce homological stratification and establish its basic properties. We begin with the following homological variant of [BHS23d, Theorem 4.1].
@@
by [BCH+23b, (13.4)]. This proves that $\operatorname{Locid}\langle f_{*}(\mathbb{1})\rangle$ is minimal. ∎
@@
Let $\mathscr{T}=\operatorname{SH}$ be the stable homotopy category. By [BAL20b, Corollary 5.10], which relies on the nilpotence theorem of Devinatz–Hopkins–Smith [DHS88], the canonical map $\operatorname{Spc}^{h}(\operatorname{SH}^{c})\to\operatorname{Spc}(\operatorname{SH}^{c})$ is a bijection; cf. 8.1 below. In particular, for a given prime number $p$, there is a homological prime $\mathscr{B}_{p,\infty}$ corresponding to the tt-prime of all finite $p$-torsion spectra. The associated weak ring identifies with the mod $p$ Eilenberg–Mac Lane spectrum $H{\mathbb{F}_{\hskip-0.81949ptp}}$, i.e., $E_{\mathscr{B}_{p,\infty}}=H{\mathbb{F}_{\hskip-0.81949ptp}}$; see [BC21, Corollary 3.6]. However, it follows from Ravenel’s work [RAV84] that there are strict inclusions
$0\subsetneq\operatorname{Locid}\langle I\rangle\subsetneq\operatorname{Locid}\langle H{\mathbb{F}_{\hskip-0.81949ptp}}\rangle$
where $I$ denotes the $p$-local Brown–Comenetz dual of the sphere spectrum. Indeed, on the one hand, we have $I\in\operatorname{Locid}\langle H{\mathbb{F}_{\hskip-0.81949ptp}}\rangle$ because the homotopy groups of $I$ are $p$-torsion and concentrated in non-negative degrees. On the other hand, we have $H{\mathbb{F}_{\hskip-0.81949ptp}}\not\in\operatorname{Locid}\langle I\rangle$ because $H{\mathbb{F}_{\hskip-0.81949ptp}}\otimes H{\mathbb{F}_{\hskip-0.81949ptp}}\neq 0$ while $H{\mathbb{F}_{\hskip-0.81949ptp}}\otimes I=0$.
@@
The spectrum in 3.14 is a single point, which shows that there is no obvious condition on the space $\operatorname{Spc}^{h}(\mathscr{T}^{c})$ that ensures that the homological local-to-global principle holds. This should be contrasted with the classical local-to-global principle, which holds whenever $\operatorname{Spc}(\mathscr{T}^{c}$) is noetherian [BHS23d, Theorem 3.22].
@@
Recall that the Bousfield class $A(t)$ of an object $t\in\mathscr{T}$ is the localizing ideal $A(t)\coloneqq\ker(-\otimes t)$ and that the Bousfield lattice of $\mathscr{T}$ is the lattice of Bousfield classes ordered by reverse inclusion; see [BHS23d, Section 8]. Surprisingly, the following proposition shows that the classification of Bousfield classes is equivalent to the h-detection property whenever the homological support agrees with the naive homological support. Indeed, it is immediate from the definition of naive homological support that $\operatorname{Supp}^{n}(t)\subseteq\operatorname{Supp}^{n}(s)$ whenever $A(t)\leq A(s)$. Therefore, there is a well-defined order-preserving map
$\big\{\text{Bousfield classes of }\mathscr{T}\big\}\rightarrow\big\{\text{subsets of }\operatorname{Spc}^{h}(\mathscr{T}^{c})\big\}$
@@
For any $\mathscr{T}$ satisfying the h-detection property, we obtain a description of the meet operation on its Bousfield lattice: $A(s)\wedge A(t)=A(s\otimes t)$; cf. [BHS23d, Theorem 8.8].
@@
If $\mathscr{T}$ is h-stratified then every localizing ideal is a Bousfield class and the lattice of localizing ideals of $\mathscr{T}$ is isomorphic to the Bousfield lattice of $\mathscr{T}$; cf. [BHS23d, Theorem 8.8].
@@ 4. Homological stratification via cosupport
We now establish characterizations of homological stratification in terms of the behaviour of homological cosupport. The proof of the next proposition closely follows its tensor-triangular counterpart ([BCH+23b, Theorem 7.15]):
@@
so $\mathsf{hom}(t_{1},t_{2})\neq 0$ by $(c)$. The minimality criterion [BCH+23b, Lemma 7.12] then shows that $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle$ is minimal. ∎
@@
4.2 is a homological analogue of the characterization of stratification in terms of cosupport [BCH+23b, Theorem 7.15]. The goal of the rest of this section is to establish a variant of this result formulated purely in terms of support and cosupport.
@@
The proof of the next theorem makes use of Brown–Comenetz duality. Recall from [BCH+23b, Definition 9.4] that the Brown–Comenetz dual $I_{c}$ of a compact object $c\in\mathscr{T}$ is characterized by the formula
@@
for all $t\in\mathscr{T}$. Following the proof of [BCH+23b, Proposition 12.9], we compute